OpenStudy (samsan9):

Convert this y=x^2-2x-3 into vertex form, and intercept form

OpenStudy (aum):

Complete the square of x^2-2x-3

OpenStudy (samsan9):

i got (x-3) (x+1)

OpenStudy (aum):

Not factor. Complete the square. x^2 - 2x - 3 = (x - 1)^2 - 1 - 3 = (x - 1)^2 - 4 y = (x - 1)^2 - 4 is the vertex form.

OpenStudy (aum):

But for the intercept form, what you did is correct. y = (x - 3)(x + 1) is the intercept form.

OpenStudy (samsan9):

so the intercept form is always factored ?

OpenStudy (aum):

Yes, for a quadratic equation, the intercept form is the factored form. y = (x - 3)(x + 1) is called the intercept form because looking at the factors you can tell the x-intercepts. When x = 3 or -1, y becomes 0. And that is the definition of x-intercepts: What value of x will make y zero? When plotted those are the x values where the graph will cross the x-axis.

OpenStudy (samsan9):

what about standard form?

OpenStudy (samsan9):

is it just making it into an ax^2+bx+c?

OpenStudy (aum):

Yes. The equation given in the problem is already in the standard form: y = x^2 - 2x - 3.

OpenStudy (samsan9):

how do you convert a vertex into an intercept form?

OpenStudy (aum):

If you expand the vertex form and simplify, you will get the standard form. If you factor the standard form you will get the intercept form.

OpenStudy (samsan9):

thank you

OpenStudy (aum):

You are welcome.